OpenStudy (anonymous):

How would I solve x+3 = 2x-3? (Their both under a radical sign)

hero (hero):

square both sides first

OpenStudy (anonymous):

square both sides then you get x+3=2x-3 subtract x on both sides 3=2x-3-x 3=x-3 add three on both side 3+3=x or x=6

OpenStudy (anonymous):

Kay I got that

hero (hero):

radical sign and square cancels because they are inveses of each other